OpenStudy (anonymous):

What is the scale factor of ABC to XYZ? A. 5 B. 3 C. D. one third

OpenStudy (anonymous):

A. 5 B. 3 C. D. one third

OpenStudy (alexandervonhumboldt2):

divide any side of XYZ by corresponding side of ABC. for example YX/BA=9/45=1/5